7. What is the mass of :
(a) 1 mole of nitrogen atoms?
(b) 4 moles of aluminium atoms (atomic mass of aluminium = 27)?
(c) 10 moles of sodium sulphite (Na2SO3)?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) The mass of 1 mole of nitrogen atoms is 14 g. (b) The mass of 4 moles of aluminium atoms is (4 × 27) g = 108 g (c) The mass of 10 moles of sodium sulphite (Na SO ) is 10 × [2 × 23 + 32 + 3 × 16) g = 10 × 126 g = 1260 g